# AI-driven View Guidance System in Intra-cardiac Echocardiography Imaging

### v1

@@ 3.3 Dataset
To achieve our goal, we need to model a transition function $\mathcal{M}$. For this, we utilize a dataset of ICE images that are labeled according to the anatomical structures they visualize. These images are paired with corresponding tip position and orientation, which are obtained from the electromagnetic (EM) sensors, providing precise tracking of the catheter’s spatial configuration. Ultimately, the dataset consists of pairs of images with their corresponding anatomical labels and position/orientation information. Additionally, the corresponding volume mesh is required to obtain the target information. For this purpose, we reconstruct volumetric information of cardiac structures using volume contouring algorithms [36]. It generates intra-cardiac images targeting the LA, LAA, Left Inferior Pulmonary Vein (LIPV), Left Superior Pulmonary Vein (LSPV), Right Inferior Pulmonary Vein (RIPV), Right Superior Pulmonary Vein (RSPV), and ESO, as shown in Figure 3 (a). Each dataset contains a target volume mesh with world coordinate system-based positioning, several images visualizing each cardiac part, and the corresponding position and orientation information. We used a dataset collected from the CARTO system (Biosense Webster Inc., USA) [37], comprising data from 858 subjects, with 793 used for training, 25 for validation, and 40 for testing.
3.4 Missing mesh
Our dataset mainly consisted of the LA, LAA, LIPV, LSPV, RIPV, and RSPV volumes. Thus, it includes only a limited dataset for the RA, RV, and LV volumes, comprising five subjects. However, the RV and LV are crucial for intra-cardiac examinations, with the RV view often serving as the home view—the starting point for most procedures.
To overcome this limitation, we estimated the centers of the RA, RV, and LV volumes based on other structures, such as the center points of the LA and LAA, using a subject dataset that includes RA, RV, and LV volumes. This allowed us to determine the relationships between these centers. By applying this relationship to our whole dataset, we estimated the missing centers of the RV, LV, and RA, as illustrated in Figure 3 (b).
@@ 3.5 Data pre-processing
In the dataset, each subject has its own coordinate system, known as the world coordinate system, which can vary due to the EM sensors used during data acquisition, leading to inconsistencies across subjects. To address this, we normalized all datasets based on the transducers used in each image, eliminating dependence on the world coordinate system. This normalization is also more practical since transducer manipulation should be performed relative to the transducer itself.
The transducer state in the world coordinate system can be represented as ${\bf S}_{w}^{tr}$ and its corresponding transformation matrix is $T_{w}^{tr}$. We can obtain the target position and orientation relative to the transducer following below equation:
@@
where the ${\bf T}_{w}^{tar}$ (${\bf S}_{w}^{tar}$) is target position and orientation, which visualizing each center of the mesh. According to clinical guidance in ICE imaging procedures, the structures are visualized by rotating the ICE, with minimal variation in position. Therefore, we selected a single point inside the RA volume for each target view as target position. To determine the orientation, we aligned the fan direction (target position to center point of fan long axis) to point toward the center of the mesh. The final positions and orientations, ${\bf S}_{tr}^{tar}$, was utilized our target dataset. The composition of the entire dataset for a single subject is shown in Figure 3 (c), while the individual targets for a single subject and single view are shown in Figure 3 (d).
4 Implementation Details
@@ 4.1 Network details
We utilized MedMamba as the base model for our method [30]. MedMamba combines convolutional neural networks (CNN) with Mamba, which uses State Space Models (SSM). This combination leverages both local and global information, showing superior performance in medical classification tasks compared to other conventional methods. Consequently, we applied MedMamba to our view navigation research, addressing the regression problem.
The entire architecture is shown in Figure 4. The ICE image, representing the current view, is fed into the network and processed through ResNet50 to extract features [38]. Simultaneously, the target view class, represented as a one-hot vector of size 6 (corresponding to the number of target views), is passed through a single linear layer that adjusts its length to match the squared feature size of the ResNet50 output. This is then resized and concatenated with the image features before being passed through another feature mixing layer composed of a single convolution layer.
MedMamba consists of four sequences of SS-Conv-SSM blocks and patch merging steps, except for the last block, which contains only an SS-Conv-SSM block. The SS-Conv-SSM block splits its feature map into two branches: one for local feature extraction using convolution and the other for global feature extraction using the core element, 2D-Selective-Scan (SS2D). It employs the Cross-Scan Module (CSM), which uses a four-way scanning strategy to achieve a global receptive field without increasing the computational burden. The selective scan state-space model (S6) captures long sequence dependencies from all directional features and merges them to construct the 2D feature map. The patch merging step reduces the spatial dimension to achieve hierarchical representations. At the output of MedMamba, we added two single linear layer heads: one for position regression and the other for orientation regression.
The entire network was trained using quantile loss to estimate the boundaries of estimation. This approach enables the network to estimate not only the median value but also the $q^{th}$ percentile, providing the boundary of estimation with $q^{th}$ reliability. The simplified loss function is shown below:
@@
where the $p,o$ represent the labels for position and orientation, respectively and $\hat{p},\hat{o}$ denote the predicted position and orientation, respectively. The $\lambda$ is weighting parameter.
Figure 5: Validation method. (a) Accuracy calculation: The prediction is considered correct if the predicted view matches the target view, and incorrect otherwise. (b) Distance calculation: The distance is measured from the center of the volume to the predicted view. The normalized distance is calculated by dividing distance with a half distance.
Figure 6: The qualitative results. (a) Randomly selected 12 samples which from specific view visualizing specific structure. The light gray is current view and black is 50 quantile prediction view. Each of dark gray denote the 98, 2 quantile prediction view that represents the boundaries of prediction. (b) Randomly selected 4 samples which from undefined view visualizing non-specific structure.
@@ 4.2 Evaluation metric
To evaluate the proposed method, we aimed to determine how accurately the relative position and orientation can guide the fan to the target mesh. We assessed this in two ways. First, we used classification to measure how well the fan’s position and orientation, as predicted by the network, align with the target volume. As shown in Figure 5 (a), a correct view is defined as the fan being properly positioned within the target volume, while a view where the fan is outside the target volume is considered incorrect. We calculated the accuracy of this classification.
Second, we measured the distance between the proposed fan position and the center of the target volume. Specifically, we computed two types of distance factors. As shown in Figure 5 (b), the first type is the real distance, which measures the vertical distance between the center of the target volume and the proposed fan position. The second type is the normalized distance, which standardizes the real distance based on half the size of the volume, where a distance of 1 represents the edge of the volume from its center.
@@ 4.3 Training details
The network was trained using the Adam optimizer with a learning rate of 1e-4 during 580 epochs. We employed a batch size of 40, and the entire image dataset was resized to $224\times 224$. In Figure 4, the $N$ is set to [2,2,4,2] for each block, which is default setting of MedMamba-Tiny. The $\lambda$ in Equation 10 is 10 and the $\alpha$ is 0.98 which denotes the $98^{th}$ quantile.The entire framework was implemented in PyTorch and trained on a single A100 GPU.
@@
Figure 7: Quantitative results. The x-axis is distance axis and y-axis is percentage of samples. (a) Total sample distance distribution. (b) Correct sample distance distribution. (c) Correct sample real distance distribution for each target view.
5 Experimental Results
Since the proposed method is designed for real clinical settings and trained on real clinical data, our ability to perform actual validation is limited. Therefore, we present simulation-based results derived from real clinical data to demonstrate the method’s effectiveness. To validate the method, we conducted evaluations using both qualitative and quantitative approaches. The current view image, along with each of the six target view classes, is input into the proposed network, which adjusts the fan’s position and direction based on the prediction values. In the qualitative evaluation, we visually assessed whether the predicted fan was well-contained within the target volume. For the quantitative evaluation, as detailed in Section 4.2, we calculated the success rate percentage of how well the predicted fan was included within the target volume. Additionally, we measured the distance from the center of the target volume to the predicted fan, where a smaller distance indicates a more accurate estimation.
We test our method using 40 subjects which does not overlapped with training subjects. In particular, each subject has several frames, therefore, we test all frames whether it can give desired position and orientation well depending on target view. Finally, 6532 test cases were utilized as final testset.
5.1 Qualitative result
@@
The qualitative results are shown Figure 6. From a current view, we determined the relative position and orientation focusing on a specific view among six different target views. The light gray color indicates the current view position and orientation, while black and dark gray represent the 50 and 98 or 2 percentiles of the predicted fan, respectively.
@@
As shown in the Figure 6 (a), the results include 12 randomly selected cases where the input image corresponds to specific views such as RV, LV, LPV, RPV, LAA, or ESO. In all cases, the predicted fan is clearly contained within the target volume. Particularly, for cases requiring significant movement, such as from LV to RPV, the predictions accurately reflect the desired state. Conversely, for cases requiring minimal movement, such as from LPV to LPV, the predictions also effectively facilitate view navigation. In the Figure 6 (b), it shows 4 randomly selected cases where the input image visualizes undefined views that only intersect the LA volume. In these cases, the proposed method provides accurate target state information for both small and large movements.
5.2 Quantitative result
[…]
@@ 5.2 Quantitative result
The Figure 7 (a) shows the normalized distance between the center of the target volume and the predicted fan. The histogram illustrates the distribution of these distances, including out-of-boundary (OOB) samples. The normalized distance is calculated by dividing the distance from the predicted fan to the center of the target volume by the distance from the center plane to the edge plane. Thus, a distance of 1 represents the edge of the volume, distances below 1 indicate correct predictions, and distances above 1 indicate incorrect predictions. In Figure 7 (b), it shows the histogram for correct sample distances only, with smaller distances corresponding to better predictions. In Figure 7 (c), it gives the correct sample distance distribution for each target view, visualizing histograms for various input images corresponding to target views such as RV, LV, LPV, RPV, LAA, or ESO. For the RV and LV target views, the histograms are slightly noisier due to the smaller number of training samples compared to other target views.
@@
Figure 8: Nearby sample result. ’View A’ and ’View B’ show similar or nearby view of the same area. (a) Both images show LV view (b) Both images show ESO view (c) Both images show LPV view. (d) ’View A’ shows undefined view and ’View B’ shows ESO view. (e) Both images show RPV view.
Table 1: Nearby view prediction difference results
@@
5.3 Nearby view result
Since our proposed model is trained based on real clinical data, we evaluated it through a simulation-based process. To assess the robustness of the method, we tested it with views that are similar or nearby to the target viewpoint. Specifically, the method should produce similar results or only minor differences when the current view covers a similar or adjacent area to the target viewpoint. We tested the method using 5 randomly selected samples from different subject that visualized similar or nearby areas.
@@ 5.3 Nearby view result
In the Figure 8, it shows examples of similar or nearby views. In each sample, ’View A’ and ’View B’ visualize very similar areas but from different image. We obtained target states for 6 target views from our proposed method using both images and calculated the differences. The results are summarized in Table 1.
@@
In the case of Sample 1, both views visualize the LV. When targeting the RV viewpoint, the difference between the outputs of the two views is $5.72[mm]$ and $(0.045,0.121,0.088)[rad]$ that can be demonstrated that very similar result. In most of cases, the results show a good result in position and orientation difference for each target view. We demonstrate that the proposed method can account for difference in image to the difference in coordinate.
5.4 State estimation result

### v3

@@ 3.3 Dataset
To achieve our goal, we need to model a transition function $\mathcal{M}$. For this, we used a dataset collected from a well-established clinical environment and processed using the CARTO mapping system (Biosense Webster Inc., USA) [34]. The CARTO system is an application designed for real-time visualization and mapping of cardiac anatomy. It uses an ICE catheter to scan the entire cardiac structure and reconstruct it into a 3-D volumetric mesh using volume contouring algorithms [37]. It targets the LA, LAA, Left Inferior Pulmonary Vein (LIPV), Left Superior Pulmonary Vein (LSPV), Right Inferior Pulmonary Vein (RIPV), Right Superior Pulmonary Vein (RSPV), and ESO, as shown in Figure 3 (a). Notably, it captures cardiac images at the end of the T-wave, providing consistent volumetric information. The CARTO images are labeled according to the anatomical structures they visualize. These images are paired with the corresponding tip position and orientation, which are obtained from the electromagnetic (EM) sensor. The EM sensor provides precise tracking of the catheter’s spatial configuration. Consequently, the CARTO system integrates four key elements: US images, anatomical labels, position/orientation information, and the subject’s cardiac mesh—all of which were utilized in this research. We obtained the CARTO dataset from a total of 858 subjects. Of these, 793 subjects were allocated for training, 25 for validation, and 40 for testing. This resulted in approximately 143,400 cases for training, 4,870 cases for validation, and 6,532 cases for testing.
3.4 Reconstruction of Missing Volume Meshes
Our dataset mainly consisted of LA, LAA, LIPV, LSPV, RIPV, and RSPV volumes. Thus, it contains only a limited number of samples for the RA, RV, and LV volumes, comprising five subjects. However, the RV and LV are crucial for intra-cardiac examinations, with the RV view often serving as the home view-the starting point for most procedures.
To address this limitation, we estimated the centers of the RA, RV, and LV volumes based on their relationships with other structures. Specifically, we determined the spatial relationships between each of the RA, RV or LV and the LA and LAA. By extracting these relationships from a subset of the dataset containing all these volumes, we applied the derived relationships to the entire dataset as shown in Figure 3 (b).
@@ 3.5 Data pre-processing
In the dataset, each subject has its own coordinate system, referred to as the world coordinate system. This can vary due to the EM sensors used during data acquisition, leading to inconsistencies across subjects. To address this, we normalized all datasets based on the transducers used in each image, eliminating dependence on the world coordinate system. This normalization is also more practical, as transducer manipulation should be performed relative to the transducer itself.
The transducer state in the world coordinate system can be represented as ${\bf S}_{w}^{tr}$, and its corresponding transformation matrix is $T_{w}^{tr}$. We can obtain the target position and orientation relative to the transducer following equation:
@@
where the ${\bf T}_{w}^{tar}$ (${\bf S}_{w}^{tar}$) is target position and orientation, which visualizing each center of the mesh. According to clinical guidance in ICE imaging procedures, the structures are visualized by rotating the ICE catheter with minimal variation in position. Therefore, we selected a single point inside the RA volume for each target view as the target position. To determine the orientation, we aligned the fan direction (from the target position to the center point of the fan’s long axis) to point toward the center of the mesh. The final positions and orientations, ${\bf S}_{tr}^{tar}$, were used in our target dataset. The composition of the entire dataset for a single subject is shown in Figure 3 (c), while the individual targets for a single subject and view are shown in Figure 3 (d).
@@ 3.6 Network details
We utilized MedMamba as the base model for our method [30]. MedMamba combines CNN with Mamba, which uses SSM. This combination leverages both local and global information, showing superior performance in medical classification tasks compared to other conventional methods. Consequently, we applied MedMamba to our view navigation research to solve the regression problem.
The entire MedMamba architecture is shown in Figure 4. It consists of four sequences of SS-Conv-SSM blocks and patch merging steps, except for the last block, which contains only an SS-Conv-SSM block. The SS-Conv-SSM block splits its feature map into two branches: one for local feature extraction using convolution and the other for global feature extraction using the core element, 2D-Selective-Scan (SS2D). It employs the Cross-Scan Module (CSM), which utilizes a four-way scanning strategy to achieve a global receptive field without increasing the computational burden. The selective scan state-space model (S6) captures long-sequence dependencies from all directional features and merges them to construct the 2D feature map. The patch merging step reduces the spatial dimension to achieve hierarchical representations.
To achieve our goal, we modified both the network architecture and the training loss. The original MedMamba was designed for medical image classification, which required only a single image as input. However, our objective was to develop a multi-target regression model that not only takes an image as input but also incorporates the target class. To facilitate this, we introduced a feature-mixing layer, consisting of a single convolutional layer. Specifically, the ICE image, representing the current view, is fed into the network and processed through ResNet50 to extract features [38]. Simultaneously, the target view class, encoded as a one-hot vector of size 6 (corresponding to the number of target views), is passed through a linear layer that adjusts its dimensions to match the squared feature size of the ResNet50 output. This vector is then resized and concatenated with the image features before being passed through the feature-mixing layer. Additionally, to estimate both position and orientation effectively, we added two linear layer heads at the output of MedMamba: one for position regression and the other for orientation regression.
The entire network was trained using quantile loss to account for uncertainty and variability, enabling it to effectively estimate the boundaries of predictions. This approach allows the network to estimate not only the median value but also the $q^{th}$ percentile, providing the boundary of estimation with $q^{th}$ reliability. The simplified loss function is shown below:
@@
where $p$ and $o$ represent the labels for position and orientation, respectively, and $\hat{p}$ and $\hat{o}$ denote the predicted position and orientation, respectively. The parameter $\lambda$ is the weighting factor.
4 System Details and Evaluation Metrics
Figure 5: Validation method. (a) Percentage of correct views calculation: The prediction is considered correct if the predicted view matches the target volume, and incorrect otherwise. (b) Distance calculation: The distance is measured from the center of the target volume to the predicted view.
Figure 6: Random view to clinically defined view results. The light gray is current view, and the black denotes the $50^{th}$ quantile prediction view. Each of dark gray denote the $98^{th}$, $2^{nd}$ quantile prediction views, representing the boundaries of the prediction. The orange arrow indicates the transition from the current view to the predicted view, based on the output of the proposed network. (a) Results using 12 randomly selected samples from the anatomically labeled test set, each visualizing a specific structure. (b) Results using 4 randomly selected samples from the test set labeled as undefined views, each visualizing non-specific structures only crossing the LA.
@@ 4.1 Evaluation metric
To evaluate the proposed method, we aimed to determine how accurately the relative position and orientation could guide the fan to the target mesh. We assessed this in two ways. First, we used classification to measure how well the fan’s position and orientation, as predicted by the network, align with the target volume. As shown in Figure 5 (a), a correct view is defined as the fan being properly positioned within the target volume, while a view where the fan is outside the target volume is considered incorrect. We then calculated the percentage of correct view across all test cases.
Secondly, we measured the distance between the proposed fan position and the center of the target volume. As shown in Figure 5 (b), we calculated the vertical distance between the center of the target volume and the predicted fan position.
@@ 4.2 Training details
The network was trained using the Adam optimizer with a learning rate of 1e-4 for 580 epochs. We employed a batch size of 40, and the entire image dataset was resized to $224\times 224$. In Figure 4, $N$ is set to [2,2,4,2] for each block, which is the default setting of MedMamba-Tiny. The $\lambda$ in Equation 10 is 10, and $\alpha$ is 0.98, which denotes the $98^{th}$ quantile. The entire framework was implemented in PyTorch and trained on a single A100 GPU.
4.3 Guidance system frame rate
The proposed method provides continuous guidance within a closed-loop system, ensuring timely updates for effective navigation. To evaluate its performance in this regard, we measured the inference time using a single A100 GPU, achieving a frequency of 50 [fps].
@@ 4.3 Guidance system frame rate
Figure 7: Histogram of the distance between the center of the target mesh and the predicted fan. The x-axis represents the distance, and the y-axis represents the percentage of samples. (a) Real distance distribution for all samples. (b) Real distance distribution for each target view.
5 Experimental Results and Analysis
Since the proposed method is specifically designed for real clinical applications and trained on actual clinical data, direct validation in real-world scenarios is limited. Therefore, we primarily present simulation-based results derived from the real clinical CARTO dataset, evaluated both qualitatively and quantitatively. These simulation-based evaluations assess the method’s performance by comparing predicted values with the corresponding volume mesh. For example, we adjusted the fan’s position and orientation based on the predicted values and verified their alignment with the associated volume mesh. Several simulation experiments were conducted, as shown below:
  1. A
    Random view to clinically-defined view: This experiment aims to verify whether the position and orientation provided by the proposed method effectively guide the user to the desired target view.
  2. B
    Distance error between prediction and ground truth: This experiment aims to verify whether the model has been well-trained on the training dataset by evaluating the accuracy of the predicted position and orientation compared to the ground truth.
  3. C
    Distance error between the center of the target mesh and the predicted fan: This experiment evaluates whether the predicted fan accurately intersects the target mesh by measuring its proximity to the center of the mesh.
  4. D
    Distance error from a nearby view to a clinically-defined view: This experiment assesses whether the proposed method provides consistent and similar prediction values when the current views are located nearby.
  5. E
    Distance error for intermediate views while moving towards the target view: This experiment verifies whether the proposed method can effectively guide the user to the target mesh in dynamic scenarios.
To further validate the effectiveness of the proposed method in a realistic scenario, we conducted a semi-simulation experiment as described below:
  1. F
    Human-in-the-loop validation: This experiment assesses whether the proposed method can be effectively applied in realistic scenarios, allowing engineers to directly manipulate the ICE catheter based on the guidance provided.
The entire validation procedure was simulated using MATLAB 2022a, including visualization and validation of the sensor data with the CARTO dataset. Specific details of the simulation environment are provided in each section.
We evaluated our method using 40 subjects, none of whom were part of the training dataset. Each subject included several frames, resulting in a total of 6,532 test cases used in the final test set.
5.1 Random view to clinically-defined view
The dataset we used includes anatomical labels corresponding to each ultrasound image. Among the 6,532 test cases, there are several cases labeled as RV, LV, LAA, LPV, RPV, ESO, and undefined views, with the undefined view only crossing the LA and not any other structure. We randomly selected 12 cases—two for each view—and 4 cases for the undefined view.
For each selected case, the proposed method provided the relative position and orientation required to achieve the target view. Based on these predicted states, we adjusted the fan’s current position and orientation to validate whether the predicted fan accurately crossed the target mesh.
@@ 5.1 Random view to clinically-defined view
The qualitative results are shown in Figure 6. The light gray color indicates the current view fan, including position and orientation information, while black and dark gray represent the $50^{th}$ and $98^{th}$ or $2^{nd}$ percentiles of the predicted fan, respectively.
@@
In all cases, the predicted fan is clearly contained within the target volume. Notably, for cases requiring significant movements, such as transitioning from the LV to the RPV, the predictions accurately align with the desired target state. Similarly, for cases involving minimal movement, such as navigating from the LPV to the LPV, the predictions effectively support precise view navigation. In Figure 6 (b), which demonstrates the results of transitions from undefined views to clinically-defined views, the proposed method successfully provides accurate target state information for both small and large movements.
5.2 Distance error between prediction and ground truth
We calculated the errors between the predicted values, such as position and orientation, and their corresponding ground truth. Predictions were obtained for all 6,532 test cases, and the errors were computed accordingly. The average position error was 9.75 [mm], with a standard deviation of 8.50 [mm]. For orientation, the average error was (3.48, 1.13, -1.87) [deg], with a standard deviation of (47.27, 22.24, 36.51) [deg]. These results demonstrate that the proposed network has been trained effectively, as it accurately predicts the target position and orientation based on the current view image. The relatively high average orientation error is attributed to cases where the target and current positions are close, such as the RV and LV. In such cases, the target’s large size allows for a wider range of acceptable orientations. This results in a high standard deviation in orientation; however, as demonstrated in Section 5.3, the method maintains high percentage of correct views despite this variability.
5.3 Distance error between the center of the target mesh and predicted fan
We used the evaluation method described in Section 4.1. Based on the predicted position and orientation, we adjusted the current view fan to align with the predicted fan. The proposed method also provides prediction boundaries, and we evaluated the percentage of corrects views by assessing how well the predicted quantile planes were contained within the target volume. Furthermore, we measured the distance between the predicted $50^{th}$ quantile planes and the center of the target mesh, with smaller distances indicating better predictions.
The proposed method achieved an 89%, demonstrating its strong performance in effectively guiding views.
@@ 5.3 Distance error between the center of the target mesh and predicted fan
Figure 7 (a) shows the real distance histogram for the entire test dataset, including out-of-boundary (OOB) samples. In Figure 7 (b), it gives the correct sample distance distribution for each target view, visualized as histograms for various input images corresponding to target views such as RV, LV, LPV, RPV, LAA, or ESO. For the RV and LV target views, the histograms are slightly noisier due to the smaller number of training samples compared to other target views. However, consistent with the analysis that smaller distance correlate with more accurate predictions, this pattern is evident in all cases.
@@
Figure 8: Distance error from a nearby view to clinically-defined view. ’View A’ and ’View B’ show similar or nearby view of the same area. (a) Both images show the LV view (b) Both images show the ESO view (c) Both images show the LPV view. (d) ’View A’ shows an undefined view, and ’View B’ shows the ESO view. (e) Both images show the RPV view.
Table 1: Distance error from a nearby view to clinically-defined view
@@
  • The upper value represents the difference in predicted position between the first and second views, while the lower value represents the difference in orientation between them. Each value in parentheses indicates the difference along the x-, y-, and z-axes.
5.4 Distance error from a nearby view to clinically-defined view
We randomly selected two views from five samples in different subjects, each visualizing similar or nearby areas. For each sample, we selected a target view and obtained the prediction values for each nearby view. We then calculated the error between the prediction values, with the results shown in Figure 8 and Table 1.
@@ 5.4 Distance error from a nearby view to clinically-defined view
Figure 8 shows examples of nearby or similar views. In each sample, ’View A’ and ’View B’ represent images of very similar areas but captured from different perspectives. Using our proposed method, we obtained the predicted target states for six target views from both images and computed the differences between them. The summarized results are presented in Table 1.
@@
In the case of Sample 1, both views visualize the LV. When targeting the RV viewpoint, the difference between the outputs of the two views is $5.72[mm]$ and $(0.045,0.121,0.088)[rad]$, demonstrating a very similar result. In most cases, the results show a good match in position and orientation differences for each target view. These findings demonstrate that the proposed method effectively accounts for the difference in images and translates it into differences in coordinates.
5.5 Distance error for intermediate view while moving to the target view